Let A be a diagonalizable n × n matrix and let P be an invertible n × n matrix such that B = P−1AP is the diagonal form of A. Prove that Ak = PBkP−1, where k is a positive integer. Use the result above to find the indicated power of A. A = −4 0 4 −3 −1 4 −6 0 6 , A5
Answer #1

1 solution We have A be diggenalizable nan matrix such that B=PAP where P cu an invertible nan matix Pre multiply both sidesAssume that result w true far kam 1.e. AM PBM pl we will prove the result for komt Par Kamal Amt AMA (Pomp-11.A [by (2) (P8pie la = PBkpl Now, we have 1-4 As -3 -1 4 4 -6 6 The eigen value of a are given by IA-dl=0 4 4 1-4-0 -3 -6 o =0 6-d 20 20 >>A =) d= 0,-1,2 24 Let x= X2 be the eigen vector cosoresponding to do- 93 Then (A +1 )x=0 -3 3 - 6 4 4 7 X₂ Applying & Be A an24 NE 1: w the basis of eigen sbare Quociated coich ds-1 Agein let X= be dhe egen veetcon cosetes pending to dzo *3 Then AX=0Applying Rg Rg + + 8R and Rg Ry -R, we get 1 1 0 13 4 - 1320 and g- X3 20 =) x = x3 and 2 =23 x = x 2 X3 ng 1) - 8 16. in the4 Again, let X= be the elgonveces comesponding Ma 33 00 d=2 Then (A-21 )X =0 -6 o 24 = ) - 3 -3 L-6 INI o Applying Rg + Pg -8 쁜 213 2/3 t *3 1 {Bly in the basis of eigenspace corresponding to d=2 Then we have three .L.P. eigen vectors so that A i di6 [com 0 0 P 0 o Ipt 0 0 Q35 O 0 0 = CHE P 0 1 - 0 py LO o 32 NBLO P= =) pia 3 0 -२ 0 3 0 o O a) AS o -1 o 3 0-2 2 3 o 32 0 0